on a road map with a scale of 1/4 inch per 10 miles, the highway from Waukee to Winterset is 1 3/8 inches long. How may miles long is this highway?
thank you very much for helping me i have tried but i have forgoten how to work out this problem.

Answer by stanbon(48522)

(1 3/8)/(1/4)
=(11/8)/(1/4)
=(11/8)*(4/1)
=11/2
So, 11/2 is the number of times 1 3/8 contains 1/4.
Multiply 11/2 by 10 to get the number of miles you seek.
11/2 * 10 = 55 miles
Cheers,
Stan H.